# Detectors for Safe and Reliable LLMs: Implementations, Uses, and Limitations

### v2

@@ 4.2. Why is this important?
Because social harms are the product of context-dependent classification systems with deep historical roots and are socially and morally charged, we need to pay careful attention to the choices we make in constructing the detectors. By deploying or embedding these detectors in real world applications, we are contributing to and enforcing classification systems that impose a certain order, in turn impacting human interactions and social structures (Bowker and Star, 1999).
4.2.1. Reproduction, enforcement, and scaling of harmful context and practices
@@
Since annotation means inscribing values and categorizing extracts of text, and considering that the definition of stigma is context-dependent and fluid, through annotating the stigma dataset or evaluating the detector, we might reproduce harmful stereotypes, unfair discrimination, and exclusionary norms or stigmatizing practices. If the detector is eventually integrated into IBM’s commercial platform or the dataset is open-sourced, this problematic reproduction could be scaled upwards and outwards in ways that are not easily seen or controlled.
4.2.2. Lower Performance, Usefulness, or Explainability
@@
There may be worse performance for certain social groups that have different definitions of stigma or lower performance in relation to the deployment application (the context of use), questions which still need to be investigated. When we annotate the stigma dataset based only on one person’s or culture’s perspectives, there is a high risk of neglecting not only the social, cultural, and temporal context of the data but also inadvertently neglecting the context of use (i.e., the place where model is being deployed or the output the end-user intends to mitigate). Explainability of detector results in different use cases might be significantly reduced because of the challenges mentioned and the overlapping assumptions and gaps that might occur during data annotation. Given the potential and expected difficulty disentangling various stigma considerations without additional context, applications that assign a set of text as being high in stigma might have no way to easily explain or justify this outcome in light of end-user disagreement. Returning to Figure 3, it would be challenging for the model to adequately and reliably address all or most of these questions. In absence of this rationale, how would users understand or trust its response?
@@
5. Additional Future Directions
@@
Much of the current research discussion has centered on single-turn interactions, i.e., analyzing a model’s response for a given prompt. However, as language models become more sophisticated, so does their ability to maintain a coherent dialogue over multiple turns. Detection in this multi-turn setting becomes crucial for a complete understanding of the LLM’s risk. One current focus is the development of detectors for harmful conversations. Prior work focused on detecting egregiously bad conversations between humans and non-LLM conversational agents, using key features such as repeated utterances (by the human or agent), emotional indicators, or explicitly asking for a human to detect when the conversation is turning bad (Sandbank et al., 2018; Weisz et al., 2019). When evaluating interactions between humans and LLM-driven agents it becomes necessary, given their increased sophistication, to be more careful about the (potentially subtle) ways in which conversations can turn bad. To this end, current work is focused on building detectors based on carefully designed principles of effective human-AI communication, paying particular attention to how the conversational context influences the harmfulness of a particular response.
@@
Current efforts to better understand jail-breaking attacks highlight the need for a more unified and effective detection strategy. While some attempts have been made to characterize prompt attacks (Shen et al., 2023a; Wei et al., 2023; Zeng et al., 2024), there is currently no overarching strategy for effectively detecting such attacks. Existing methods involve leveraging metrics like perplexity as features for detection (Jain et al., 2023; Alon and Kamfonas, 2023), particularly in suffix-style attacks (Zou et al., 2023), or by robust aggregation of model responses based on multiple perturbed input queries (Kumar et al., 2023; Robey et al., 2023). Additionally, moderation policies have been employed to identify natural language prompt injections (Rebedea et al., 2023). Current work is focused on expanding upon these approaches by leveraging a red-teaming pipeline, in turn laying the groundwork for a comprehensive detection strategy.
Multilinguality
It is important to support languages beyond English to meet diverse requirements. Unfortunately, non-English data is more scarce than English data, which makes creating multilingual detectors difficult. The explicit-hate detector was trained with multilinguality in mind and we will continue to expand the scope of other detectors in future work.
Acknowledgements.
@@
Appendix A Modes of Detection
Figure 4. Various detector modes. In the single-turn setting, detectors can either monitor the (a) prompt, (b) response, or (c) the prompt and response. The multi-turn setting (d) describes monitoring of a given response subject to the context provided by the history of prompts and past responses.
@@ Appendix A Modes of Detection
Appendix B Synthetic Data Generation for the Stigma Detector
As mentioned in 2.1, we used a synthetic data generation approach where we leveraged LLMs, prompted using an in-context learning style, to generate more data based on stigmas found in (Pachankis et al., 2018).
@@ Appendix B Synthetic Data Generation for the Stigma Detector
Appendix C Uncertainty Calibration and OOD Performance
As mentioned in 2.4, we added large quantities of negative (i.e., benign) labeled data. Specifically, we added the data used to train the blocklisting detectors (Mitchell, 1999) as it was readily available, legally permissible, and deemed appropriate for this task - due to the fact that the data was in the style of news headlines that did not contain any explicit content.
Initially, we saw a performance of 0.15 accuracy on this data (with around 5000 examples in the test set) (Mitchell, 1999). Note that all data points are labeled negative (i.e., “not hate”), implying that our false positive rate was 0.85. However, once we added the additional data to the fine-tuning method used to train the detector, we were able to achieve an accuracy of 0.95. Although it remains to be seen if the updated detector is over-fitting to this new data, this is still a step in the right direction, as the new data represents out of distribution examples, which the detector is more likely to see once deployed.
@@ Appendix C Uncertainty Calibration and OOD Performance
Additionally, we report results with the conformal predictor. Note that these results are for the implicit-hate detector, using the validation and test sets from the original implicit hate dataset (ElSherief et al., 2021). Both sets contain around 4000 samples, whereas the training set contained just over 12,000 samples. Our desired coverage was 90% and we achieved an empirical coverage of 90.4%. Next, 38% of data instances were abstained on. Finally, we report some metrics on the entire test set compared with the non-abstained dataset below:
@@
Appendix D User Interface
Figure 5. Red Teaming + Guardrails UI
[…]
A user interface which encourages interactive probing of both generative models and the detectors themselves. More details in 2
@@ Appendix D User Interface
Appendix E Evaluating detectors on real-world data
@@
We observed that the Mistral-7b model (Jiang et al., 2023) exhibits a favorable combination of these two criteria. Moreover, in the course of generating these questions, we observed that exercising better control over the generation process enables the creation of more refined and diversified queries, as described in (Kour et al., 2023). To create a diverse set of questions related to social stigma, we adopted a problem-breakdown approach. This involved curating a list of various stigmatized groups and their associated stigmas in a specific region (e.g., USA). Subsequently, we instructed the model to generate questions related to that minority group and specific stigma using the following prompt template:

### v3

@@ 4.2 Why is this important?
Because social harms are the product of context-dependent classification systems with deep historical roots and are socially and morally charged, we need to pay careful attention to the choices we make in constructing the detectors. By deploying or embedding these detectors in real world applications, we are contributing to and enforcing classification systems that impose a certain order, in turn impacting human interactions and social structures (Bowker and Star 1999).
Reproduction, enforcement, and scaling of harmful context and practices
@@
Since annotation means inscribing values and categorizing extracts of text, and considering that the definition of stigma is context-dependent and fluid, through annotating the dataset or evaluating the detector, we might reproduce harmful stereotypes, unfair discrimination, and exclusionary norms or stigmatizing practices. If the detector is eventually integrated into IBM’s commercial platform or the dataset is open-sourced, this problematic reproduction could be scaled upwards and outwards in ways that are not easily seen or controlled.
Lower Performance, Usefulness, or Explainability
@@
There may be worse performance for certain social groups that have different definitions of stigma or lower performance in relation to the deployment application (the context of use). When we annotate the stigma dataset based only on one person’s or culture’s perspectives, there is a high risk of neglecting not only the social, cultural, and temporal context of the data but also inadvertently neglecting the context of use (i.e., the place where model is being deployed or the output the end-user intends to mitigate).
@@
5 Additional Future Directions
@@
Much of the current research discussion has centered on single-turn interactions, i.e., analyzing a model’s response for a given prompt. However, as language models become more sophisticated, so does their ability to maintain a coherent dialogue over multiple turns. Prior work focused on detecting egregiously bad conversations between humans and non-LLM conversational agents, using key features such as repeated utterances (by the human or agent), emotional indicators, or explicitly asking for a human to detect when the conversation is turning bad (Sandbank et al. 2018; Weisz et al. 2019). When evaluating interactions between humans and LLM-driven agents it becomes necessary, given their increased sophistication, to be more careful about the potentially subtle ways in which conversations can degrade. To this end, current work is focused on building detectors based on carefully designed principles of effective human-AI communication, paying particular attention to how the conversational context influences the harmfulness of a particular response (Miehling et al. 2024).
@@
Current efforts to better understand jail-breaking attacks highlight the need for a more unified and effective strategy. While some attempts have been made to characterize prompt attacks (Shen et al. 2023b; Wei, Haghtalab, and Steinhardt 2023; Zeng et al. 2024), there is currently no overarching strategy for effectively detecting such attacks. Existing methods involve leveraging metrics like perplexity as features for detection (Jain et al. 2023; Alon and Kamfonas 2023), particularly in suffix-style attacks (Zou et al. 2023), or by robust aggregation of model responses based on multiple perturbed input queries (Kumar et al. 2023; Robey et al. 2023). Additionally, moderation policies have been employed to identify natural language prompt injections (Rebedea et al. 2023). Current work is focused on expanding these approaches by leveraging a red-teaming pipeline, in turn laying the groundwork for comprehensive detection.
Attribution
Algorithmic explanations of the detector scores can help users better understand detector behavior and provide feedback. Since the detectors are text classifiers, it is possible to use existing explanation methods (Ribeiro, Singh, and Guestrin 2016; Lundberg and Lee 2017; Sundararajan, Taly, and Yan 2017; Chen, Zheng, and Ji 2020; Kim et al. 2020; Mosca et al. 2022) to associate importance scores with spans of text, which indicate their contribution to the detector score and can be displayed by highlighting text. One challenge however is the length of the input to the detector, which may be a paragraph-length response as in Figure 2 or even longer if the detector considers the input to the LLM. Future work improves such explanation methods for long input text in terms of both computational cost and interpretability of the attributed text spans (Paes et al. 2024).
6 Acknowledgments
@@
Appendix A Implementation Details
We provide detailed information (including training data, model, and evaluation) regarding two of our detectors - the implicit hate and faithfulness detectors, below.
A.1 Implicit-Hate-Detector
In order to train the implicit-hate-detector, we used a combination of 4 datasets. We started with the Latent Hatred dataset (ElSherief et al. 2021b), which is a benchmark that was specifically designed for implicit hate speech. Then, to combat the issue of high false positives (which we elaborate on in the answer to your second question below), we use the 20 NewsGroups dataset (Mitchell 1999) - which was primarily used to train the blocklisting detectors. Note that we deliberately use this dataset in the hopes of “increasing the proportion of negative (i.e., benign) labeled data in our training set” (Section 2.4 Reliable Uncertainties). Third, we use a dataset from a work titled Identifying Implicitly Abusive Remarks about Identity Groups using a Linguistically Informed Approach (this is the Identity Groups row in the table below) (Wiegand, Eder, and Ruppenhofer 2022). Finally, we add in a subset of the CivilComments dataset (Borkan et al. 2019), taking only samples which have an identity_attack column value of greater than 0.5, which we believe corresponded to implicitly hateful comments.
As is the case with most of our detectors, we took the uncased BERT model from HuggingFace (specific model link here: https://huggingface.co/google-bert/bert-base-uncased). During training, we use a batch size of 16, we start with a learning rate of 0.000001, and we train for 50 epochs, taking the best model with respect to validation f1 score.
For evaluation, please refer to Table 1.
test dataset accuracy balanced accuracy Precision Recall F1
implicit-hate 0.754 0.747 0.616 0.724 0.665
blocklisting 0.676 0.676 - - -
identity groups 0.752 0.732 0.729 0.891 0.802
civil comments 0.974 0.974 1.0 0.974 0.987
Table 1: Evaluation for implicit-hate-detector
We note that:
  • blocklisting data only contains benign (i.e. negatively or 0-labeled examples). hence, precision/recall/f1 do not apply (they are trivially equal to 0)
  • When evaluating, we predominantly focus on f1 score, in order to balance both false positives and false negatives. However, from the point of view of an end user, we would argue that a false negative is more egregious and harmful. A false negative indicates that a piece of harmful text is classified as benign, thus potentially displaying harmful text to the end user - this applies when the detector is used in the “guardrail” modality.
A.2 Faithfulness-Detector
For the faithulness-detector, we used the Multi-Genre Natural Language Inference (MultiNLI) (Williams, Nangia, and Bowman 2018) and the Stanford Natural Language Inference (SNLI) (Bowman et al. 2015) datasets. Additionally, we also generated around 22.5k synthetic data MRQA datasets (we used HotPotQA (Yang et al. 2018) and NewsQA (Trischler et al. 2017)) which we mixed with the above two datasets.
For the model, we finetuned a deberta-v3-large model on the three above datasets using binary NLI labels.
For evaluation, we compute the ROC-AUC of these models for a variety of datasets in Table 2. Our model shows better ROC-AUC numbers with respect to metrics of comparable model sizes (BertScore, BARTScore, FactCC). Note that the ANLI metric uses an 11B T5 model that is much larger than deberta model.
v1 v2 v2_mix BertScore BARTScore FactCC ANLI (11B)
FRANK 84 89.0 86.7 84.3 86.1 76.4 89.4
SummEval 69.4 81.4 78.3 77.2 73.5 75.9 80.5
MNBM 73.2 53.2 75.1 62.8 60.9 59.4 77.9
QAGS-C 82.5 88.2 86.9 69.1 80.9 76.4 82.1
QAGS-X 73.8 73.7 79.9 49.5 53.8 64.9 83.8
BEGIN 76.5 48.7 79.1 87.9 86.3 64.4 82.6
Q2 74.1 82.5 77.9 70 64.9 63.7 72.7
DialFact 84.1 76.2 89.2 64.2 65.6 55.3 77.7
PAWS 80.5 80.4 86.6 77.5 77.5 64 86.4
Avg 77.6 74.8 82.2 71.4 72.2 66.7 81.5
Table 2: Evaluation for faithfulness-detector
Some notes:
  • v1 refers to: Our model (Deberta-v3) + MNLI/SNLI only
  • v2 refers to: Our model v2 (Deberta-v3) + Mixtral synthetic data
  • v2_mix refers to: Our model v2 (Deberta-v3) + Mixtral synthetic data + MNLI/SNLI
Appendix B Modes of Detection
Figure 4: Various detector modes. In the single-turn setting, detectors can either monitor the (a) prompt, (b) response, or (c) the prompt and response. The multi-turn setting (d) describes monitoring of a given response subject to the context provided by the history of prompts and past responses.
@@ Appendix B Modes of Detection
Appendix C Synthetic Data Generation for the Stigma Detector
As mentioned in 2.1, we used a synthetic data generation approach where we leveraged LLMs, prompted using an in-context learning style, to generate more data based on stigmas found in (Pachankis et al. 2018).
@@ Appendix C Synthetic Data Generation for the Stigma Detector
Appendix D Uncertainty Calibration and OOD Performance
As mentioned in 2.4, we added large quantities of negative (i.e., benign) labeled data. Specifically, we added the data used to train the blocklisting detectors (Mitchell 1999) as it was readily available, legally permissible, and deemed appropriate for this task - due to the fact that the data was in the style of news headlines that did not contain any explicit content.
Initially, we saw a performance of 0.15 accuracy on this data (with around 5000 examples in the test set) (Mitchell 1999). Note that all data points are labeled negative (i.e., “not hate”), implying that our false positive rate was 0.85. However, once we added the additional data to the fine-tuning method used to train the detector, we were able to achieve an accuracy of 0.95. Although it remains to be seen if the updated detector is over-fitting to this new data, this is still a step in the right direction, as the new data represents out of distribution examples, which the detector is more likely to see once deployed.
@@ Appendix D Uncertainty Calibration and OOD Performance
Additionally, we report results with the conformal predictor. Note that these results are for the implicit-hate detector, using the validation and test sets from the original implicit hate dataset (ElSherief et al. 2021a). Both sets contain around 4000 samples, whereas the training set contained just over 12,000 samples. Our desired coverage was 90% and we achieved an empirical coverage of 90.4%. Next, 38% of data instances were abstained on. Finally, we report some metrics on the entire test set compared with the non-abstained dataset below:
@@
Appendix E User Interface
Figure 5: Red Teaming + Guardrails UI: A user interface which encourages interactive probing of both generative models and the detectors themselves. More details in 2
@@ Appendix E User Interface
Appendix F Evaluating detectors on real-world data
@@
We observed that the Mistral-7b model (Jiang et al. 2023) exhibits a favorable combination of these two criteria. Moreover, in the course of generating these questions, we observed that exercising better control over the generation process enables the creation of more refined and diversified queries, as described in (Kour et al. 2023). To create a diverse set of questions related to social stigma, we adopted a problem-breakdown approach. This involved curating a list of various stigmatized groups and their associated stigmas in a specific region (e.g., USA). Subsequently, we instructed the model to generate questions related to that minority group and specific stigma using the following prompt template: